In 24 years, Travis will be 3 times his present his present age. How old is he now?

Answer by crazylogms(2)

lets presume, the present age of travis to be :- X years
after 24 years his age becames :- X+24 years
when he attends the age of X+24 years he becomes trice of it's present age (3X years)
Therefore,

X+24 = 3X
Doing some algebra we get

2X = 24
dividing each side by 2

X = 12
Thus the present age of Travis is 12 years.
After 24 years his age is ( 12 + 24 ) = 36 years which is exactly trice of it's present age ( 3 * 12 = 36 years )
Cheers.
Crazyloms !!!!!!!!